Please, I need help.
for the line 2x-5y=852
find the slope?
find the slope as a line to it
thanks sooooo much

Answer by PRMath(133)

Please, I need help.
for the line 2x-5y=852
find the slope?
find the slope as a line to it
thanks sooooo much

Hi...


Don't panic - you can do this. The equation you are given: 2x - 5y = 852 just needs to be put into a different format. You need to put the line into the
y = mx + b equation of a line, which is a SLOPE/INTERCEPT form of an equation of a line.

In the y = mx + b format, the y intercept is always "b" and the slope is always represented by "m". :-)


How do we put 2x - 5y = 852 into the y = mx + b format? You just have to solve for y. Sooo, let's try it:

2x - 5y = 852 (subtract 2x from both sides)
-5y = - 2x + 852 (now divide -5 from both sides)
y = 2/5x - 70 and 2/5

Can you see that 2/5 is your slope? (And I don't mean the 2/5 that came after the -70. I mean you should see the slope as the 2/5 that is next to the "x" in the equation. The 2/5 figure is the "m" of your equation (and remember "m" represents the slope.)

You may wonder why you have a POSITIVE 2/5? Well, you had a -2 that was divided by a -5. Therefore, a negative divided by a negative is a POSITIVE.

SO - your slope is 2/5, and slope is always rise over run.
Hope this helps. :-)